OpenStudy (anonymous):

as x approaches 0, (x^2 + 2x)/ x^4. how do i find the limit of the function algebraically?

OpenStudy (zzr0ck3r):

divide top and bottom by x^4 and what do you get?

OpenStudy (anonymous):

idk.........

OpenStudy (zzr0ck3r):

do you know what divide means?

OpenStudy (anonymous):

yes... how do i divide a larger exponent into a smaller one

OpenStudy (zzr0ck3r):

x^2/x^4 = 1/x^2

OpenStudy (zzr0ck3r):

so you will have ((1/x^2)+(2/x^3)) / 1 does this make sense?

OpenStudy (zzr0ck3r):

you might need to go back and study your algebra before you do calculus, you should know how to divide exponents with like bases if you are going to do calc

OpenStudy (anonymous):

does the limit not exist?

OpenStudy (zzr0ck3r):

I simply divided every term by x^4 x^2/x^4 = 1/x^2 2x/x^4 = 2/x^3 x^4/x^4 = 1

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

because now we have 1/0+2/0 -> infinity